If the economy booms, Meyer&Co. stock will have a return of 18.4 percent. If the economy goes into a recession, the stock will have a loss of 7.5 percent. The probability of a boom is 72 percent while the probability of a recession is 28 percent. What is the standard deviation of the returns on the stock?

11.63%

14.75%

13.52%

9.69%

15.89%

15.89%
Answer #1

Expected return=Respective return*Respective probabilities

=(18.4*0.72)+(-7.5*0.28)=11.148

Probability Return Probability*(Return-Mean)^2
0.72 18.4 0.72*(18.4-11.148)^2=37.86588288
0.28 -7.5 0.28(-7.5-11.148)^2=97.36941312
Total=135.235296

SD=[Total Probability*(Return-Mean)^2/Total Probability]^(1/2)

=11.63%(Approx).